The ages of a man and his son add up to 39years.In 3 years the man will b 4 times as old as his son.Find the ages of the man and his son.
Found 2 solutions by oberobic, richwmiller:
Answer by oberobic(2304)

x = man's age now
y = son's age now
.
x + y = 39
so
x = 39-y
.
In three years:
x+3 = man's future age
y+3 = son's future age
.
The man's age will be 4 times his son's age:
(x+3) = 4(y+3)
.
Substitute
.
(39-y+3) = 4(y+3)
.
Solve
.
42-y = 4y + 12
30-y = 4y
30 = 5y
5y = 30
y = 6
.
x+y = 39
x + 6 = 39
x = 33
.
Check: In three years, will the man's age be 4 times the son's age?
x+3 = 36
y+3 = 9
x+3 = 4(y+3)
Correct.
.
Answer: The man is 33, and the son is 6.
.
Done.

Answer by richwmiller(17219)

m+s=39
m+3=4*(s+3)
m=33
s=6
check
33+3=4*(6+3)
36=4*9
36=36
ok